OpenStudy (anonymous):

Simplify 2x(x^2-16)/x(16-x^2)

OpenStudy (anonymous):

\[\frac{x^2-16}{16-x^2}=-1\] is a start

OpenStudy (earthcitizen):

\[2 (16-x^2) (x^2-16)\]

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

the answer -2

OpenStudy (earthcitizen):

lol, prolly

OpenStudy (kropot72):

Yes -2 is the correct answer. First factorise the terms in numerator and denominator that are the difference of two squares:\[(2x(x-4)(x+4))\div(x(4-x)(x+4))\]

OpenStudy (kropot72):

The previous expression simplifies to: \[(2x ^{2}-8x)\div(-x ^{2}+4)=(2(x ^{2}-4))\div(-x ^{2}+4)\] Multipling numerator and denominator by -1 gives: \[(-2(-x ^{2}+4))\div((-x ^{2}+4)=-2\]